Week 30 July 21 - 25


Completed tasks
--------------------

BL/WA Photographed first batch of 82 political posters in special archives.
      Photography continues. 

MB The new version of Al Mashriq is up and running, updated the Green Line things
   installed the necessary available viewers for downloading somewhere.
   Indexing and searching is working.
   Commented out parts that are irrelevant or don't work here.
   Some cleaning up left, and broken pictures left to download.

BL/WA Sample project on Solar Disinfection of Drinking Water.
   Scanned (OCR) english text papers.

PA Adding information to Ras Beirut satellite images.

PA Campus tour training QTVRs continued
   There is a problem with either the camera or the CCD being out of
   vertical in the SONY training camera that is creating slanted
   QTVR panoramas. 

MB VRML2 player for MAC downloaded from SGI

BL Sample page of digitized images of Gross Specimens from the Pathology Lab
   done.

BL Worked on standard graphics for pages.

   We have persistent problems with TCI/IP (?) on the Mac workstation, linux
   server and the DDC4 PC workstation. The Mac as the same software installed as
   my laptop which is consistently problem free with regard to the net. New OS
   7.5.3 (from an unmarked CD delivered by the Apple supplier) and OT 1.1.2
   were installed with no better result. UTP wire was replaced and the machine
   came back online. The Mac has been going problem free for about 2 weeks and
   then gives up suddenly.

   Hub was replaced with another, but the problems remained the same.

   The problem with the linux (ddc) server is that there is a 10-fold difference
   in upload (not download) speeds depending on wat machine is used.
   ddc1 (Mac laptop) to ddc : slow
   ddc2 (IIvi) to ddc : fast
   ddc3 (Mac 8600) to ddc : fast
   ddc4 (PC) to ddc : slow

   We have decided to live with the problem until the new machines
   arrive (or not). If they do not arrive next week, we will first
   change the ethernet card in the linux server (dd). Maher has also
   promised to test the wiring, so that we can acertain that both
   hub and wiring are OK if the Apple dealer has to be called in fr support.

Transferrable to week 32
------------------------

BL/PA Adobe has requested a correct Illustrated version of the
      Lebanese flag.

BL Start planning format of technical guidlines for ddc.

WA Contact Armand Krikorian (Beep 896 or voice paging) for Medical
   Students Home Page.

WA/BL Copyright work continues

BL Continue background information on guidelines.

BL/WA Sample project on Solar Disinfection of Drinking Water.
   Make a proposal for Dr. Acra's approval.
   Scan (OCR) english text papers.
   Investigate copyright releases from original publishers (UNICEF and IDRC)
   Dr. Acra has also suggested doing a project on Lebanese amber fossils.

MB Install and document usage of QTVR2 player for Windows.

PA Quided tour QTVR training continues. Atempt to adjust vertical
   alignment of camcorder.

-------------
970714/970728 bl

(Patrick Abisalloum, Wissam Adib, Marwan Badran, Barre Ludvigsen)